# ModernBERT or DeBERTaV3? Examining Architecture and Data Influence on Transformer Encoder Models Performance

### v1

@@ Abstract
Pretrained transformer-encoder models like DeBERTaV3 and ModernBERT introduce architectural advancements aimed at improving efficiency and performance. Although the authors of ModernBERT report improved performance over DeBERTaV3 on several benchmarks, the lack of disclosed training data and the absence of comparisons using a shared dataset make it difficult to determine whether these gains are due to architectural improvements or differences in training data. In this work, we conduct a controlled study by pretraining ModernBERT on the same dataset as CamemBERTaV2, a DeBERTaV3 French model, isolating the effect of model design. Our results show that the previous model generation remains superior in sample efficiency and overall benchmark performance, with ModernBERT’s primary advantage being faster training and inference speed. However, the new proposed model still provides meaningful architectural improvements compared to earlier models such as BERT and RoBERTa. Additionally, we observe that high-quality pre-training data accelerates convergence but does not significantly improve final performance, suggesting potential benchmark saturation. These findings show the importance of disentangling pretraining data from architectural innovations when evaluating transformer models.
@@ 1 Introduction
Despite the widespread adoption of decoder-only large language models (LLMs) in our post-ChatGPT era, encoder-only transformers such as BERT (Devlin et al., 2019) continue to play a central role in many NLP applications. These models remain the backbone for a wide range of non-generative tasks such as classification, named entity recognition (NER), and retrieval-based systems, especially in high-throughput or latency-sensitive environments. Their relatively low compute requirements and strong performance across standard benchmarks make them a practical choice for large-scale deployment, including in Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) pipelines or guardrails systems (Neill et al., 2024).
@@
This uncertainty motivates our study aimed to evaluate the impact of architecture versus training data by conducting a controlled comparison among ModernBERT, DeBERTaV3, and RoBERTa models. We select CamemBERTaV2 (Antoun et al., 2024), a French DeBERTaV3 model trained from scratch, as our primary reference since both its intermediate checkpoints and training dataset are publicly available. Additionally, we include CamemBERTv2 (Antoun et al., 2024), a RoBERTa (Liu, 2019) based model pretrained on the same dataset, to comprehensively assess how ModernBERT’s architectural advancements compare not only against the latest DeBERTaV3-based models but also against more traditional BERT/RoBERTa architectures. Leveraging these resources, we pretrained a French ModernBERT using the exact same dataset as CamemBERTaV2 and CamemBERTv2, thus ensuring that differences in performance directly reflect architectural variations rather than dataset composition or quality. Furthermore, we pretrained another ModernBERT variant on a carefully curated, high-quality French corpus to further explore the role of dataset quality in model performance.
@@
  • When controlling for dataset differences, DeBERTaV3 outperforms ModernBERT in terms of overall benchmark performance and training sample efficiency, indicating that DeBERTaV3’s architecture and training objective optimization provide superior learning capabilities compared to ModernBERT’s efficiency-oriented design.
@@
Our findings show the importance of clearly separating respective effects of architectural changes and training datasets when evaluating NLP models. Our controlled comparison using the same pretraining dataset provides more accurate insights into the strengths and limitations of ModernBERT and DeBERTaV3 architectures
To foster further research and reproducibility, we publicly release both of our pretrained French ModernBERT models, including the variant trained on the CamemBERTaV2 dataset and on our high-quality filtered corpus. These models, along with training scripts, intermediate checkpoints, and evaluation results, are available on HuggingFace [^1] https://huggingface.co/collections/almanach/moderncamembert-67f7e6d85ede5f7cfc1ce012.
2 Literature Review
@@
While the authors of ModernBERT report superior benchmark performance over DeBERTaV3, the absence of transparent training data and lack of head-to-head comparisons on shared datasets introduces ambiguity. It is thus unclear whether the reported improvements are driven by architectural enhancements or the underlying training data.
@@ 3.1 Pre-training Datasets
We first make use of the publicly available French dataset originally used by the authors of CamemBERTaV2. This dataset has 275 billion tokens, sourced from:
  • CulturaX-FR Corpus (Nguyen et al., 2023): The French subset of a multilingual corpus containing around 265 billion French tokens, constructed from multiple snapshots of OSCAR and mC4 corpora.
@@
We also feature a second significantly larger 1T tokens French dataset created by applying heuristic and semantic filters in addition to full deduplication to the French section of the RedPajamaV2 corpus (Weber et al., 2024), combined with the HALvesting corpus and French Wikipedia. Semantic filtering was done by fine-tuning a BERT classifier trained on a document quality dataset automatically labeled by LLama-3 70B (Grattafiori et al., 2024).
@@ 3.2 Model Training
We pretrained two variants of the ModernBERT model, one on the CamemBERTaV2 dataset, which we call ModernBERT-CV2, and the other on our high-quality filtered dataset, named ModernBERT-HQ, periodically saving intermediate checkpoints. [^2] We use the publicly available ModernBERT codebasehttps://github.com/AnswerDotAI/ModernBERT Both variants are base-sized models trained with Masked Language Modeling (MLM) for 1 trillion tokens, with a maximum sequence length of 1024, and use the same tokenizer as CamemBERTaV2. We maintained the rate of dynamic token masking to 30%, while retaining all other hyperparameters consistent with those of ModernBERT-base. Training was done on 48 NVidia H100 80GB GPUs using Pytorch’s FSDP full sharding with bfloat16 mixed precision to speed up training.
@@ 4.2 Downstream Results Analysis
NER QA CLS PAWS-X XNLI
Model F1 F1 EM Acc Acc Acc
@@
Table 1: Downstream tasks results. Bold indicates best score overall while underline indicates best score between the ModernBERT models. ModernBERT-CV2 is the ModernBERT model trained on the same data as CamemBERTaV2 while ModernBERT-HQ is the one trained on the high-quality filtered dataset. Scores are the 5-seed average of the best performing set of hyperparamters for each model.
Architecture Impact.
Comparing the models trained on identical datasets (ModernBERT-CV2 and CamemBERTaV2/CamemBERTv2), we observe that ModernBERT-CV2 consistently outperforms CamemBERTv2, thus showing ModernBERT’s improvements over BERT/RoBERTa. However, it fails to surpass CamemBERTaV2 on any task, even though the latter being only trained for a single epoch on the dataset compared to three epochs (1T tokens) for ModernBERT-CV2. This clearly demonstrates that while ModernBERT offers valuable throughput-driven architectural enhancements, these improvements do not match the contextual learning capabilities provided by DeBERTaV3’s disentangled attention and RTD-based pretraining objective.
Data Quality Impact.
Interestingly, switching to our high-quality filtered dataset (ModernBERT-HQ) only marginally improved performance on downstream tasks, despite the dataset containing three times more unique tokens than the original CamemBERTaV2 dataset. ModernBERT-HQ slightly outperformed ModernBERT-CV2 on QA (FQuad), CLS, and XNLI tasks, but improvements remained within small margins. This limited gain suggests two potential explanations: either current transformer architectures exhibit diminishing returns when exposed to additional data beyond a certain threshold, or standard French benchmarks are becoming saturated and unfit to measure model quality with further improvements in model performance. The latter possibility stresses the need for more challenging and diverse benchmarks that can effectively capture the improvements brought by higher-quality data.
4.3 Pre-training Dynamics and Sample Efficiency
@@
The training curves (shown in Figures 1 and  2) indicate that CamemBERTaV2 reaches higher performance significantly earlier in training compared to ModernBERT-CV2. The DeBERTaV3-based model’s faster improvement rate confirms its better sample efficiency due to optimizations like RTD and gradient-disentangled embedding sharing (GDES). Moreover, in scenarios where pre-training data is limited or scarce, its architectures might be more advantageous.
@@
When comparing ModernBERT-CV2 and ModernBERT-HQ training curves, we observed that the model trained on the higher-quality dataset achieved its performance plateau faster, indicating that improved data quality enhances training efficiency and accelerates convergence. Yet, it does not substantially increase the final task-specific performance scores, further confirming the hypothesis of saturation effects on standard NLP benchmarks.
@@ 4.4 Context Length Extension and Final Model Release
NER QA CLS PAWS-X XNLI
Model F1 F1 EM Acc Acc Acc

### v2

@@ Abstract
Pretrained transformer-encoder models like DeBERTaV3 and ModernBERT introduce architectural advancements aimed at improving efficiency and performance. Although the authors of ModernBERT report improved performance over DeBERTaV3 on several benchmarks, the lack of disclosed training data and the absence of comparisons using a shared dataset make it difficult to determine whether these gains are due to architectural improvements or differences in training data. In this work, we conduct a controlled study by pretraining ModernBERT on the same dataset as CamemBERTaV2, a DeBERTaV3 French model, isolating the effect of model design. Our results show that the previous model generation remains superior in sample efficiency and overall benchmark performance, with ModernBERT’s primary advantage being its support for long context, faster training, and inference speed. However, the new proposed model still provides meaningful architectural improvements compared to earlier models such as BERT and RoBERTa. Additionally, we observe that high-quality pre-training data accelerates convergence but does not significantly improve final performance, suggesting potential benchmark saturation. These findings show the importance of disentangling pretraining data from architectural innovations when evaluating transformer models.
@@ 1 Introduction
Despite the widespread adoption of decoder-only large language models (LLMs) in our post-ChatGPT era, encoder-only transformers such as BERT (Devlin et al., 2019) continue to play a central role in many NLP applications. These models remain the backbone of a wide range of non-generative tasks such as classification, named entity recognition (NER), and retrieval-based systems, especially in high-throughput or latency-sensitive environments. Their relatively low compute requirements and strong performance across standard information benchmarks benchmarks make them a practical choice for large-scale deployment, including in Retrieval-Augmented Generation (Lewis et al., 2020) pipelines or guardrails systems (Neill et al., 2024). Notable examples include Google’s EmbeddingGemma (Vera et al., 2025), the BGE family of 33 models Chen et al. (2024); Xiao et al. (2023), and multilingual encoders such as multi-ligual modernBERT Marone et al. (2025) and EuroBERT Boizard et al. (2025), alongside GTE-ModernBERT Li et al. (2023); Zhang et al. (2024).
@@
This uncertainty motivates our study aimed to evaluate the impact of architecture versus training data by conducting a controlled comparison among ModernBERT, DeBERTaV3, and RoBERTa models. We select CamemBERTaV2 (Antoun et al., 2024; 2023), a French DeBERTaV3 model trained from scratch, as our primary reference since both its intermediate checkpoints and training dataset are publicly available. Additionally, we include CamemBERTv2 (Antoun et al., 2023; 2024), a RoBERTa (Liu, 2019) based model pretrained on the same dataset, to comprehensively assess how ModernBERT’s architectural advancements compare not only against the latest DeBERTaV3-based models but also against more traditional BERT/RoBERTa architectures. Leveraging these resources, we pretrained a French ModernBERT using the exact same dataset as CamemBERTaV2 and CamemBERTv2, thus ensuring that differences in performance directly reflect architectural variations rather than dataset composition or quality. In addition, we pretrained another ModernBERT variant on a carefully curated, high-quality French corpus to further explore the role of dataset quality in model performance.
@@
  • When controlling for dataset differences, DeBERTaV3 outperforms ModernBERT in terms of overall benchmark performance and training sample efficiency, with the notable exception of text retrieval tasks where DeBERTaV3 fails completely. This indicates that while DeBERTaV3’s architecture and training objective optimization provide superior learning capabilities compared to ModernBERT’s efficiency-oriented design, they do not generalize to all tasks.
@@
Our findings show the importance of clearly separating respective effects of architectural changes and training datasets when evaluating NLP models. Our controlled comparison using the same pretraining dataset provides more accurate insights into the strengths and limitations of ModernBERT and DeBERTaV3 architectures.
To promote further research and ensure reproducibility, we publicly release our two pretrained French ModernBERT models, collectively named ModernCamemBERT, including one trained on the CamemBERTaV2 dataset and another on our high-quality filtered corpus. These models, along with intermediate checkpoints, and evaluation results, are available on HuggingFace [^1] https://huggingface.co/collections/almanach/moderncamembert.
2 Related Works
@@
While ModernBERT has gained popularity other models have also made significant contributions to the field. For instance, MosaicBERT Portes et al. (2023) was the first to focus on enhancing training efficiency and performance by increasing masking rates, sequence packing and FlashAttention. NomicBERT Nussbaum et al. (2025) introduced architectural improvements like SwiGLU activation functions, RoPE positional encoding and extended context lengths, enhancing its ability to handle longer sequences up to 2048 tokens. NeoBERT Breton et al. (2025) further advanced these developments by optimizing the depth-to-width ratio and doubling the context length, while also significantly increasing training corpus size.
We chose to compare against ModernBERT because it was the best available model at the moment we started our experimentation and it is the most popular encoder in the field. While the authors of ModernBERT report superior benchmark performance over DeBERTaV3, the absence of transparent training data and lack of head-to-head comparisons on shared datasets introduces ambiguity. It is thus unclear whether the reported improvements are driven by architectural enhancements or the underlying training data.
@@ 3.1 Pre-training Datasets
We first make use of the publicly available French dataset originally used by the authors of CamemBERTaV2 (Antoun et al., 2024). This dataset has 275 billion tokens, sourced from:
  • CulturaX-FR Corpus (Nguyen et al., 2023): The French subset of a multilingual corpus containing around 265 billion French tokens, constructed from multiple snapshots of OSCAR Ortiz Suárez et al. (2019); Abadji et al. (2021; 2022) and mC4 Xue et al. (2021) corpora.
@@
We also feature a second significantly larger 1T tokens French dataset created by applying heuristic and semantic filters in addition to full deduplication to the French section of the RedPajamaV2 corpus (Weber et al., 2024), combined with the HALvesting corpus and French Wikipedia. Semantic filtering was done following the FineWeb-Edu (Penedo et al., 2024) methodology. This method has been effective in increasing the overall quality of a corpus used for LLM training and has been widely adopted in the literature Li et al. (2024); Su et al. (2025). First, We annotated 200K samples from the RedPajamaV2 dataset with quality labels (low, medium, high), using the LLama-3 70B model (Grattafiori et al., 2024) and the prompt provided in Appendix A. This annotated subset was then used to fine-tune XLM-V-base (Liang et al., 2023), which we use to annotate the whole RedPajamaV2 corpus. The semantic score was combined with the perplexity score from a language model trained on Wikipedia, as included in the RedPajama dataset. The RedPajama authors categorize data into head, middle, and tail buckets based on perplexity. We only select data if it is either from the head bucket or has a high score from the quality classifier, while disgrading any tail or low labeled documents.
@@ 3.2 Model Training
We pretrained two variants of the ModernBERT model, one on the CamemBERTaV2 dataset, which we call ModernBERT-CV2, and the other on our high-quality filtered dataset, named ModernBERT-HQ, periodically saving intermediate checkpoints. [^2] We use the publicly available ModernBERT codebase https://github.com/AnswerDotAI/ModernBERT Both variants are base-sized models trained with Masked Language Modeling (MLM) for 1 trillion tokens, with a maximum sequence length of 1024, and use the same tokenizer as CamemBERTaV2. We maintained the rate of dynamic token masking to 30%, while retaining all other hyperparameters consistent with those of ModernBERT-base. Training was done on 48 NVidia H100 80GB GPUs using Pytorch’s FSDP full sharding with bfloat16 mixed precision to speed up training [^3] See pretraining hyperparameter details in Appendix B.
@@
NER QA CLS PAWS-X XNLI MTEB
Model F1 F1 EM Acc Acc Acc Avg
[…]
Table 1: Downstream tasks results. Bold indicates best score overall while underline indicates best score between the ModernBERT models. ModernBERT-CV2 is the ModernBERT model trained on the same data as CamemBERTaV2 while ModernBERT-HQ is the one trained on the high-quality filtered dataset. Scores are the 5-seed average of the best performing set of hyperparamters for each model. MTEB scores are the average over all tasks. Full MTEB scores are available in Table 6
@@ 4.1 Downstream Evaluation Tasks
  • Text Retrieval: We used the French subset of the translated Semantic Textual Similarity (STS) benchmark (May, 2021) for training and then evaluated the resulting models using the French Massive Text Embedding Benchmark (MTEB) (Ciancone et al., 2024; Enevoldsen et al., 2025; Muennighoff et al., 2022).
@@ 4.2 Downstream Results Analysis
Architecture Impact.
Comparing the models trained on identical datasets (ModernBERT-CV2 and CamemBERTaV2/CamemBERTv2), we observe that ModernBERT-CV2 consistently outperforms CamemBERTv2 with the exception of text retrieval, thus showing ModernBERT’s improvements over BERT/RoBERTa. However, it fails to surpass CamemBERTaV2 on any non-retrieval task, even though the latter being only trained for a single epoch on the dataset compared to three epochs (1T tokens) for ModernBERT-CV2. This clearly demonstrates that while ModernBERT offers valuable throughput-driven architectural enhancements, these improvements do not match the contextual learning capabilities provided by DeBERTaV3’s disentangled attention and RTD-based pretraining objective. Our results also confirm the observation that DeBERTaV3 fails on text embedding tasks. Despite its strong performance on NLU tasks, its sentence representations are poorly structured in the embedding space.
Data Quality Impact.
Interestingly, switching to our high-quality filtered dataset (ModernBERT-HQ) only marginally improved performance on downstream tasks, despite the dataset containing three times more unique tokens than the original CamemBERTaV2 dataset. ModernBERT-HQ slightly outperformed ModernBERT-CV2 on QA (FQuad), CLS, XNLI and text retrieval tasks, but improvements remained within small margins. This limited gain suggests two potential explanations: either current transformer architectures exhibit diminishing returns when exposed to additional data beyond a certain threshold, or standard French benchmarks are becoming saturated and unfit to measure model quality with further improvements in model performance. The latter possibility stresses the need for more challenging and diverse benchmarks that can effectively capture the improvements brought by higher-quality data.
4.3 Pre-training Dynamics and Sample Efficiency
NER QA CLS PAWS-X XNLI MTEB
Model F1 F1 EM Acc Acc Acc Avg
@@ 4.3 Pre-training Dynamics and Sample Efficiency
MLDR (NDCG@10)
Model Context Max Avg
[…]
Table 3: Maximum and highest 5-seed averaged NDCG@10 score for the Multi Long Doc Retrieval task.
@@
The training curves (shown in Figures 1 and  2) indicate that CamemBERTaV2 reaches higher performance significantly earlier in training compared to ModernBERT-CV2. The DeBERTaV3-based model’s faster improvement rate strongly suggests its better sample efficiency is due to optimizations like RTD and gradient-disentangled embedding sharing (GDES). Moreover, in scenarios where pre-training data is limited or scarce, its architectures might be more advantageous.
@@
When comparing ModernBERT-CV2 and ModernBERT-HQ downstream performance throughout the training on the challenging QA tasks 1, we observed that the model trained on the higher-quality dataset achieved its performance plateau faster, indicating that improved data quality enhances training efficiency and accelerates convergence. Yet, it does not substantially increase the final task-specific performance scores, further confirming the hypothesis of saturation effects on standard NLP benchmarks.
@@ 4.4 Context Length Extension and Final Model Release
Impact on Downstream Performance.
We observe in Table 2, that ModernBERT-CV2 provides modest gains in NER (+0.14 F1) and QA (+0.34 F1 / +0.53 EM), while performance slightly decreases on classification (CLS: –0.32 Acc, XNLI: –0.43 Acc), with PAWS-X remaining stable. Meanwhile, ModernBERT-HQ-final displays clear improvements in QA (+1.08 F1 / +0.59 EM) and CLS (+0.88 Acc), while maintaining stable results on PAWS-X and XNLI. Although NER and text retrieval performance drops slightly (–0.47 F1 and -0.64), the overall trend indicates that high-quality long-context pretraining primarily benefits tasks requiring deeper semantic understanding or longer-range dependencies.
To better understand the models’ long context extrapolation, we evaluate all models trained on the French STS dataset from earlier on the French test subset of the Multi Long Doc Retrieval (MLDR) and present the results in Table 3. The results clearly demonstrate the importance of the long-context pretraining stage. As expected from the MTEB evaluation, CamemBERTaV2 performs poorly on this retrieval benchmark, further highlighting the unsuitability of its architecture for sentence embedding tasks. Both ModernBERT variants show remarkable improvements after the final training phase, with context-extended ModernBERT-HQ achieving the highest max and average score.
However, the performance of ModernBERT-CV2 (the ModernBERT variant trained on the original CamemBERTaV2 dataset) is unexpectedly poor, scoring significantly lower than both CamemBERTV2 and its counterpart trained on our high-quality dataset, ModernBERT-HQ. We currently lack a definitive explanation for this gap. One hypothesis relates to the fundamental differences between the two datasets used in our study. The original CamemBERTaV2 dataset (275B tokens) consists primarily of web-crawled content from CulturaX-FR (constructed from OSCAR and mC4 snapshots). In contrast, our high-quality filtered dataset (1T tokens) underwent extensive semantic filtering using FineWeb-Edu methodology, perplexity-based selection, and full deduplication on the RedPajamaV2 corpus, resulting in a more coherent and diverse collection of texts. The original dataset’s heavy reliance on statistical filters may lack the semantic coherence necessary for learning robust long-form representations, whereas our filtered dataset’s emphasis on high-quality, diverse sources appears better suited for supporting complex semantic understanding required in retrieval tasks.
4.5 Downstream Training Stability
Furthermore, during hyperparameter tuning of the final checkpoint, we found the newer architecture to be particularly sensitive to learning rate choices.  [^4] The search space included multiple learning rates (5e-5 to 5e-4), schedulers, batch sizes, and seeds. Despite hyperparameter tuning, the instability persisted. Further investigation revealed that NaN values appear in the loss from the very first training batch. This strongly supports our hypothesis of a numerical instability in the underlying implementation (e.g., FlashAttention) rather than a simple hyperparameter mismatch.
4.6 Training Efficiency
In addition to model accuracy, training efficiency is a crucial factor in practice, impacting both resource costs and environmental footprint. For pretraining time, our ModernBERT training required 1300 H100 GPU-hours to complete one trillion tokens. On the other hand, CamemBERTv2 took roughly 2100 GPU-hours to train on the same dataset size while CamemBERTaV2 required around 2700 GPU-hours to complete just a single epoch, despite processing one-third of the tokens (275B). This clearly demonstrates ModernBERT’s efficient architecture advantages and its practical edge during training. However, it should be noted a significant portion of the speedup over DeBERTaV3-based models comes from engineering optimizations such as unpadding and FlashAttention, both of which are not implemented in the DeBERTa models at the time of this study.
5 Discussion
Confirming recent results Warner et al. (2024); Breton et al. (2025), our experiments showed a weakness in DeBERTaV3, which failed at the text retrieval task, despite its strong performance across other benchmarks. We hypothesize that its architecture may lack the mechanisms to build a global document representation suitable for retrieval using pooling of token embeddings. In contrast, ModernBERT performed well on this task, suggesting its alternating global and local attention layers provide a suitable architecture for text retrieval.
However, the strengths of ModernBERT in text retrieval did not extend to all tasks. Our experiments showed a deficiency in the ModernBERT question-answering experiments. We suspect that ModernBERT’s architecture, may face challenges in learning the long-range dependencies that are needed for QA tasks, particularly when informative signals are sparse. For instance, the model needs to propagate the signal from the correct answer start/end token across long distances to the question context, which resides at the beginning of the input. We hypothesize that while local attention layers are efficient, they might occasionally interrupt the direct propagation path for such sparse signals. The global attention layers are then tasked with bridging these segments, leading to potential difficulties in predicting the answer span based on the distant question. Other tasks such as NER, which rely on token-level prediction, do not require long-range dependencies since the model can infer the correct label based on local context, in addition to the richer learning signal, since named entity tokens are more frequent than a single start/end token per example.
Our work contributes to the renewed debate on encoder architectures versus the prevailing trend of using decoder-only models for all tasks (see Gisserot-Boukhlef et al. (2025) and references therein). This discussion centers on whether a single, large architecture can be adapted for any task, or if specialized models remain superior for certain domains like NLU. The recent paper by Weller et al. (2025) offers new evidence by training paired encoder and decoder models under the same conditions. Their experiments confirm that encoders are better suited for classification and retrieval, which aligns with our results showing ModernBERT’s strength in text retrieval. Furthermore, the paper demonstrates that adapting a model to a task for which it was not designed is an ineffective strategy. This puts our own results into a broader context, suggesting that specialized encoders are not obsolete, and that architectural choice remains a key factor for task performance.
Looking ahead, future work could explore integrating recent efficiency improvements such as FlashDeBERTa [^5] https://github.com/Knowledgator/FlashDeBERTa, which applies FlashAttention to the disentangled attention mechanism and greatly reduces DeBERTa’s memory and latency costs. Another promising direction is investigating why DeBERTaV3 fails so strongly on embedding-based tasks. Examining its pooling mechanisms and representation geometry may help clarify the limits of disentangled attention for retrieval-oriented objectives.
6 Conclusion
We set out in this work to critically evaluate the claims made in the original ModernBERT paper by reproducing its setup under tightly controlled conditions. We isolated the authors’ contributions by retraining their model under the same conditions as the previous state-of-the-art models, to assess their actual impact on training dynamics, efficiency, and downstream performance.
Our findings show that while ModernBERT does offer improvements in training and inference speed compared to older architectures, these do not translate into better sample efficiency or task performance under matched conditions. In fact, under careful evaluation, we found that DeBERTaV3’s architecture and training objectives are more advantageous in low-data scenarios or if the goal is to get the absolute best task performance, except for the case of text retrieval where DeBERTaV3 fails completely.
We also observed that increasing the size and quality of pretraining data only yielded marginal gains for the newly proposed architecture, suggesting that current benchmarks may be reaching saturation, or at least they are insufficiently sensitive to capture finer improvements. During fine-tuning, we faced a problem with sensitivity to hyperparameters, which the V2 baselines did not have. These stability concerns present challenges for reproducibility and deployment, and deserve further investigation.
Limitations
Our study has several important limitations. First, we observe that ModernBERT exhibits training instability during fine-tuning, this might be due to numerical instabitlity of the flashattention implementation. Second, our downstream evaluation relies on established NLP benchmarks that may be reaching saturation, potentially masking more nuanced performance differences between architectures. Finally, our analysis focuses on base-sized models, and the relative performance characteristics may differ for larger model variants. Future work should address these limitations through stability analysis, development of more discriminative benchmarks, and evaluation across different model scales.
Ethics Considerations
Acknowledgments
This work has received partial funding Benoît Sagot and Djamé Seddah’s chairs in the PRAIRIE-PSAI, funded by the French national agency ANR, as part of the “France 2030” strategy under the reference ANR-23-IACL-0008. This project also received funding from the BPI Scribe project. The authors extend their gratitude to the OPAL infrastructure of Université Côte d’Azur for providing essential resources and support. This work was also granted access to the HPC resources of IDRIS by GENCI under the allocation 2024-GC011015610 and AD011013900R2. Special thanks to Nathan Godey and Francis Kulumba for their assistance with training code and for engaging in productive discussions.
All pretraining dataset are sourced from publicly available and wildly used corpora (CultraX from OSCAR [CC0 1.0], and mc4 [odc-by], HAL’s open archive [CC], Wikipedia [cc-by-sa-3.0], RedPajamaV2 [Common Crawl Permissive license]). Downstream dataset are all from standard open benchmarks. ModernBERT’s codebase is Apache-2.0. CamemBERTaV2, ModernCamembert checkpoints and models are released under the MIT licence. [^6] https://huggingface.co/collections/almanach/ moderncamembert
Appendix A Quality Labeling Prompt
[…]
Appendix B Pretraining Details and Hyperparamters
We closely follow the original ModernBERT recipe. We present the model parameters in Table 4 and pretraining hyperparamters in Table  5
Parameter Base
Vocabulary 32,768
Unused Tokens 418
Layers 22
Hidden Size 768
Transformer Block Pre-Norm
Activation Function GeLU
Linear Bias False
Attention Multi-head
Attention Heads 12
Global Attention Every three layers
Local Attention Window 128
Intermediate Size 1,152
GLU Expansion 2,304
Normalization LayerNorm
Norm Epsilon 1e-5
Norm Bias False
RoPE theta 160,000
Local Attn RoPE theta 10,000
Table 4: ModernBERT model parameters
Pretraining Context Extension Context Ext & High Quality
Training Tokens 1 trillion 150 billion 100 billion
Max Sequence Length 1,024 8,192 8,192
Batch Size 4,608 768 768
Warmup (tokens) None
Microbatch Size 96 8 8
Learning Rate 8e-4 3e-4 3e-4
Schedule Trapezoidal Trapezoidal 1-sqrt (50B tokens delayed)
Warmup (tokens) 3 billion None None
Weight Decay 1e-5
Training Time (hours) 22.7 4.5 4
Model Initialization Megatron - -
Dropout (attn out) 0.1
Dropout (all other layers) 0.0
Optimizer DecoupledAdamW
Betas (0.90, 0.98)
Epsilon 1e-06
Training Hardware 48 GPUs - 12x(4xH100)
Training Strategy FSDP - Full Sharding
Software Libraries PyTorch 2.5.1, Cuda 12.4, Composer 0.28, Flash Attention 2.6.3-Hopper
Table 5: Pre-training hyperparameters
Appendix C Detailed MTEB Scores
Clustering Classification Pair Classification Retrieval Reranking STS Summarization Overall